### v1

# May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability

Abstract.
Research in explainable AI (XAI) aims to provide insights into the decision-making process of opaque AI models. To date, most XAI methods offer one-off and static explanations, which cannot cater to the diverse backgrounds and understanding levels of users. With this paper, we investigate if free-form conversations can enhance users’ comprehension of static explanations, improve acceptance and trust in the explanation methods, and facilitate human-AI collaboration. Participants are presented with static explanations, followed by a conversation with a human expert regarding the explanations. We measure the effect of the conversation on participants’ ability to choose, from three machine learning models, the most accurate one based on explanations and their self-reported comprehension, acceptance, and trust. Empirical results show that conversations significantly improve comprehension, acceptance, trust, and collaboration. Our findings highlight the importance of customized model explanations in the format of free-form conversations and provide insights for the future design of conversational explanations.
Explainable AI (XAI), Conversation, Interpretability, Interactive Explanation, Human-AI Interaction, XAI for Computer Vision
ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Computing methodologies Artificial intelligence
1. Introduction
The rapid advancement of Artificial Intelligence (AI) is powered by opaque deep neural networks (DNNs), which are difficult to interpret by humans (Bodria et al., 2023). The lack of transparency prevents verification of AI decisions by human domain experts and is especially concerning in areas of high-stake decisions, such as healthcare and law enforcement, where erroneous algorithmic decisions could lead to severe harm to human lives (Caruana et al., 2015; Cai et al., 2019) and erosion of public trust (Quinn et al., 2021; Powles and Hodson, 2017). To improve the interpretability (or explainability, which we use interchangeably in this paper) of AI models, numerous eXplainable Artificial Intelligence (XAI) methods have been proposed (see survey articles (Yang et al., 2019; Danilevsky et al., 2020; Bodria et al., 2023)). Perhaps surprisingly, despite the research effort, a number of recent studies (Poursabzi-Sangdeh et al., 2021; Zhang et al., 2020; Bansal et al., 2021; Wang and Yin, 2021; Lakkaraju et al., 2022; Liao et al., 2020) find that the explanations often do not resolve user confusion regarding the neural networks they are purported to explain. This curious phenomenon warrants further investigation.
Having reviewed the literature, we attribute user dissatisfaction to three major reasons. First, the explanations do not properly account for layperson users’ knowledge of machine learning, which may not be sufficient to mentally establish the causal relation between the explanations and the model behaviors being explained (Poursabzi-Sangdeh et al., 2021; Ma et al., 2023; Springer and Whittaker, 2019). Communication theory posits that effective communication requires the experts to establish common ground with layperson users (Clark and Marshall, 1981; Clark and Brennan, 1991). However, experts usually find it hard to accurately estimate what laypeople know (Wittwer et al., 2008; Wilkesmann and Wilkesmann, 2011; Miller, 2019). To make matters worse, underestimating and overestimating the user’s knowledge level are equally detrimental to communication (Wittwer et al., 2008; Lakkaraju et al., 2022). As a result, it is little wonder that the explanations designed by experts are almost always at a mismatch with the user’s actual knowledge level.
Second, existing techniques are designed to explain only one facet of model behaviors, and follow-up questions naturally arise (Liao et al., 2020; Lakkaraju et al., 2022; Abdul et al., 2018), but existing interaction paradigms typically do not account for this. For example, feature attribution techniques (Sundararajan et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2013; Ribeiro et al., 2016; Hu et al., 2018; Alvarez-Melis and Jaakkola, 2017) reveal the input features that the model rely on to make a prediction, but does not show how predictions would change if the input features change. This may be answered using counterfactual explanations (Verma et al., 2020; Wachter et al., 2017; Sharma et al., 2019; Karimi et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021). However, current interaction paradigms of model explanation typically supply only a single static explanation, and assumes that explanation has been accomplished after that; they are not prepared for the possibility of follow-up questions.
Finally, users of XAI have diverse backgrounds, intentions, and information needs (Suresh et al., 2021; Wang and Yin, 2021; Ehsan et al., 2021; Liao et al., 2020). For example, Liao and Varshney (Liao and Varshney, 2021) identifies five different objectives of users of explanations, including model debugging, assessing the capabilities of AI systems, making informed decisions, seeking recourse or contesting the AI, and auditing for legal or ethical compliance. Naturally, one static explanation is not sufficient to satisfy all objectives and purposes. Therefore, researchers (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020; Dazeley et al., 2021; Cheng et al., 2019) have suggested injecting interactivity to model explanations in order to establish common ground, address knowledge gaps, and to create customized explanations that adapt to the users.
Existing work on interactive explanations can be broadly categorized into two types. The first type, interactive machine learning (Amershi et al., 2014; Fails and Olsen Jr, 2003), generates model explanations that users rely on to provide feedback and suggestions to the machine learning model. Their primary goal is to improve machine learning performances of by interacting with machine learning practitioners, rather than explaining model behaviors to layperson users. In this interactive setting, explanations have been shown to improve user satisfaction (Smith-Renner et al., 2020) and feedback quality (Liang et al., 2020; Kulesza et al., 2015). The second type aims to demonstrate the input-output relationship by allowing users to freely modify input features and observe changes in the output, while showing current feature attribution explanations (Cheng et al., 2019; Liu et al., 2021a; Tenney et al., 2020; Hohman et al., 2019). This type of interactivity has been shown to improve user understanding (Cheng et al., 2019) and perceived usefulness (Liu et al., 2021a) of AI models. However, the effective use of these interactive approaches still requires a rudimentary understanding of achine learning, such as the generic relation between input and output, or what model properties the interpretations reveal. These interactive explanations still cannot answer most types of follow-up questions laypeople may have about the explanations.
Free-form conversations that accompany static explanations are arguably the most versatile mode of interaction as they allow users to ask arbitrary follow-up questions and can give tailored explanations of neural networks for users with diverse backgrounds (Liao et al., 2020; Dazeley et al., 2021; Feldhus et al., 2022; Lakkaraju et al., 2022). Through interviews with decision-makers, Lakkaraju et al. (Lakkaraju et al., 2022) discover that they have a strong preference for explanations in natural language dialogue. They also outline five principles of interactive explanations and argue that conversational explanations fulfill all five. Hence, the authors argue that conversations are an appealing form of explanation for users with limited machine learning knowledge. With the progress in conversational characters (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023), AI systems that can answer questions about their own decisions appear to be within our reach in the near future. However, before investing effort to develop such a chatbot, it is beneficial to empirically quantify the effects of conversational explanations.
In the current study, we conduct Wizard-of-Oz experiments to investigate how conversations assist users in understanding static explanations, improving acceptance and trust in XAI methods, and selecting the best AI models based on explanations. Specifically, we first present participants with examples of static explanations and measure their objective understanding and subjective perception of static explanations. Afterward, the participants carry a conversation with an AI system through a textual interface and seek to clarify any doubts. An XAI expert acts as the mock AI system and attempts to answer the questions faithfully through the textual interface. The participants are motivated to understand the explanations as they need to select the best-performing machine learning model only using explanations when doing objective evaluation. After the conversation, participants complete the same pre-conversation measurements. By comparing results before and after the conversation, we can estimate the effects of conversations on static explanations.
We design both an objective evaluation to measure users’ understanding of static explanations and a subjective evaluation to gauge users’ perceptions. The subjective evaluation can complement the objective evaluation to provide a holistic perspective on the users’ grasp of explanations (Hoffman et al., 2018; Cheng et al., 2019). In the objective evaluation, we ask users to choose, using the explanations, one neural network from three candidates that is the most accurate on unobserved data. The users are presented with several input instances, on which the three neural networks make exactly the same decisions. The only differences between the three networks lie in their explanations. Hence, to make the correct selection, the users must understand the explanations. The subjective evaluation contains 13 questions requiring users to self-report their perceptions of static explanations from three aspects: comprehension, acceptance, and trust of static explanations. The questions are adopted from existing works that have been validated in prior research.
The acceptance of explanations by end users, i.e. their intentions to use explanations, indicates the effectiveness and usability of those explanations. We measure users’ acceptance by their perceived usefulness, ease of use, and behavioral intention of explanations. These three aspects are derived from the Technology Acceptance Model (TAM) (Davis, 1989; Davis et al., 1989; Diop et al., 2019), one of the most influential and widely applied theories to explain individuals’ acceptance and use of information systems.
Trust has been recognized as an important factor in human-AI collaboration as it mediates the human’s reliance on AI models, thus directly affecting the effectiveness of the human-AI collaboration (Sebo et al., 2020; Seaborn et al., 2021; Doshi-Velez and Kim, 2017). However, the literature reports mixed effects of explanations on trust. Though several studies have shown that increased transparency in machine learning systems does not significantly impact people’s trust (Cheng et al., 2019; Schaffer et al., 2019; Kunkel et al., 2019; Poursabzi-Sangdeh et al., 2021), others find that explanations can improve users’ trust (Larasati et al., 2020; Yang et al., 2020) and can potentially lead to misplaced trust in machine learning models (Wang and Yin, 2021; Zhang et al., 2020). Given these contrasting views, we measure users’ trust in explanations to investigate the effects of conversations.
From the results, we discover that free-form conversations with XAI experts in the Wizard-of-Oz setting significantly improve participants’ comprehension of static explanations and their ability to select the best AI models based on explanations. Besides, they report an increase in perceived acceptance and trust of static explanations after conversations. By analyzing the collected conversations, we observe that participants’ questions are unanticipated and diverse. This explains why static explanations do not resolve users’ confusion about neural networks and indicates the importance and necessity of conversational explanations. We also discuss the implication of our study for the future design of dialogue systems to explain static explanations. Our study underscores the effects of free-form conversations on neural network explainability in practice and provides insights into the future development of conversational explanations. To the best of our knowledge, we present the first study that investigates the effects of Wizard-of-Oz conversations on neural network explainability in practice.
2. Related Work
Explainable Artificial Intelligence (XAI) refers to those models that can explain either the learning process or the outcome of AI predictions to human users (Yang et al., 2019). In this section, we will discuss related work about static and interactive explanation methods in XAI.
2.1. Static Explanation
Static XAI is usually categorized into two groups: self-explanatory models and post-hoc methods. Post-hoc methods can be categorized into feature attribution methods and example-based methods. Self-explanatory models are inherently transparent, offering clarity in their decision-making processes and facilitating explainability (Danilevsky et al., 2020; Bodria et al., 2023). Self-explanatory models can be viewed as the opposite of black box models. Examples of such models include linear regression, logistic regression, decision trees, Naive Bayes, attention mechanism (Bahdanau et al., 2014), decision sets (Lakkaraju et al., 2016), rule-based models (Rudziński, 2016; Yang et al., 2017), among others. However, the requirements of self-explanatory models place constraints on model design, which may cause them to underperform in complex tasks. Conversely, the majority of recent XAI methods are post-hoc XAI methods, which can be used for an already developed or trained model that is usually not inherently transparent (Selvaraju et al., 2017; Ribeiro et al., 2016; Chen et al., 2021; Verma et al., 2020; Adadi and Berrada, 2018; Bodria et al., 2023). These methods often do not attempt to explain how the model works internally, but instead, employ separate techniques to extract explanatory information. Post-hoc XAI methods can be viewed as reverse engineering processes that employ other independent explanatory models or techniques to extract explanatory information without altering, elucidating, or even understanding the inner workings of the original black-box model. There are two main groups of methods to generate post-hoc XAI explanations, i.e., feature attribution methods and example-based methods.
2.1.1. Feature Attribution Methods
Feature attribution methods (Sundararajan et al., 2017; Selvaraju et al., 2017; Cortez and Embrechts, 2013; Simonyan et al., 2013; Lundberg and Lee, 2017; Ribeiro et al., 2016; Hu et al., 2018; Alvarez-Melis and Jaakkola, 2017; Liu et al., 2018; Shih et al., 2018; Ignatiev et al., 2019) explain model predictions by investigating the importance of different input features to final predictions. There are two main types of feature attribution methods, gradient-based methods (Cortez and Embrechts, 2013; Sundararajan et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2013; Lundberg and Lee, 2017) and surrogate methods (Ribeiro et al., 2016; Hu et al., 2018; Alvarez-Melis and Jaakkola, 2017; Liu et al., 2018; Shih et al., 2018; Ignatiev et al., 2019). Gradient-based methods use gradients/derivatives to evaluate the contribution of a model input on the model output. An example method is Grad-CAM(Selvaraju et al., 2017). It superimposes a heatmap on the regions of important input features by weighting the activations of the final convolutional layer by their corresponding gradients and averaging the resulting weights spatially. Besides directly calculating the importance score of input features, several methods propose to use a simple and understandable surrogate model, e.g., a linear model, to locally approximate the complex deep neural model. Surrogate models can explain the predictions from the complex deep neural model due to their inherent interpretable nature. LIME and its variants are typical methods for generating local surrogate models. LIME (Ribeiro et al., 2016) builds a linear model locally around the data point to be interpreted and generates an instance-level explanation for the output.
2.1.2. Example-based Methods.
Example-based methods (Chen et al., 2021; Verma et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021; Jeyakumar et al., 2020) refer to those that explain predictions of black-box models by identifying and presenting a selection of similar or representative instances. Those examples can be selected or generated from different perspectives, such as training data points that are the most influential to the parameters of a prediction model or the predictions themselves (Chen et al., 2019, 2021; Yoon et al., 2020), counterfactual examples that are similar to the input query but with different predictions (Verma et al., 2020; Wachter et al., 2017; Sharma et al., 2019; Karimi et al., 2020; Tran et al., 2021; Mothilal et al., 2020; Poyiadzi et al., 2020; Wijaya, 2021), or prototypes that contain semantically similar parts to input instances (Croce et al., 2019; Bien and Tibshirani, 2011; Jeyakumar et al., 2020; Mikolov et al., 2013; Doshi-Velez et al., 2015; Kim et al., 2016).
In this work, we mainly focus on feature attribution methods as they directly highlight the importance of input features, making the model’s decision-making process more intuitive (Kim et al., 2023). Specifically, we select Grad-CAM from gradient-based methods and LIME from surrogate methods to conduct conversational explanations with participants.
2.2. Interactive Explanation
Several studies emphasize the need for interactivity in XAI methods (Lakkaraju et al., 2022; Abdul et al., 2018; Schmid and Wrede, 2022; Rohlfing et al., 2020). For instance, Lakkaraju et al. (Lakkaraju et al., 2022) find that decision-makers strongly prefer interactive explanations. Similarly, a literature analysis by Abdul et al. (Abdul et al., 2018) suggests that interactions can help users progressively explore and gather insights from static explanations. Rohlfing et al. (Rohlfing et al., 2020) reason that explanations should be co-constructed in an interaction between the explainer and the explainee, adapting to individual differences since the human understanding process is dynamic. From an interdisciplinary perspective, Schmid and Wrede (Schmid and Wrede, 2022) underscore the necessity of user-XAI interactions to adapt to diverse information requirements.
To integrate interactivity and explainability, two primary methodologies emerge. One group of methods focuses on using explanations to help users provide feedback about improving machine learning models. In these methods, the interactivity lies in the cycle of model explanation, user feedback, and model improvement. Explanations aim to help users better understand model decisions and provide valuable feedback. As a result, machine learning models can be incrementally trained with additional loss terms from explanatory feedback (Kulesza et al., 2015; Lertvittayakumjorn et al., 2020; Liang et al., 2020; Ross et al., 2017; Schramowski et al., 2020; Smith-Renner et al., 2020) or with added data points (Alkan et al., 2022; Biswas and Parikh, 2013; Teso and Kersting, 2019; Teso et al., 2021). However, these methods are aimed at machine learning practitioners who can well understand and utilize explanations. Another group focuses on enhancing user understanding of explanations by allowing them to modify the model input and observe changes in the corresponding output. Such interactivity has been shown to improve user comprehension and the perceived utility of AI models (Cheng et al., 2019; Liu et al., 2021a). For instance, Tenney et al. (Tenney et al., 2020) and Hohman et al. (Hohman et al., 2019) propose different user interfaces that allow for debugging and understanding machine learning models by examining input-output relationships. However, a rudimentary understanding of machine learning is still required for effective utilization of these interfaces, such as the generic relation between input and output, or what model properties the interpretations reveal.
HCI researchers have recently proposed that XAI methods should align with the ways humans naturally explain mechanisms. Specifically, Lombrozo (Lombrozo, 2006) argues that an explanation is a byproduct of a conversational interaction process between an explainer and an explainee. Miller (Miller, 2019) argues that explanations should contain a communication process, where the explainer interactively provides the information required for the explainee to understand the causes of the event through conversations. Hilton(Hilton, 1996) suggests a two-stage process: diagnosis to identify causation and then an explanation to communicate the diagnosis. Building on this perspective of human explanations, recent works envision ”explainability as dialogue” to provide explanations suitable for a wide range of layperson users (Liao et al., 2020; Dazeley et al., 2021; Feldhus et al., 2022; Lakkaraju et al., 2022). While there is much theoretical analysis about the significance of conversations in explainability, practical investigations into their impact on users remain limited. In this study, we explore the role of conversations in enhancing users’ comprehension of static explanations, and how they affect users’ acceptance, trust, and collaboration with these explanations.
2.3. Human-AI Collaboration
Human-AI collaboration is an emerging research area, which explores how humans and AI systems can work together to achieve shared goals (Arous et al., 2020; Lai et al., 2022; Kim et al., 2023). Prior studies within this domain have investigated collaborations between humans and various AI systems, from robots (Gero et al., 2020; Häuslschmid et al., 2017; Kumar et al., 2021; Ogenyi et al., 2021) to virtual agents (Bickmore et al., 2009; Ashktorab et al., 2020; Cai et al., 2019; D’Avella et al., 2022; Numata et al., 2020). The tasks involved span a broad scope, including text (Bansal et al., 2021) and image (Kim et al., 2023) classifications, medical diagnosis (Cai et al., 2019), deception detection (Lai and Tan, 2019) and cooperative games (Gero et al., 2020; Ashktorab et al., 2020; Feng and Boyd-Graber, 2019). An area of particular interest within these collaborations is the role of explanations in influencing human-AI decision-making (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021).
Our study aligns with existing work on human-AI collaboration (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021; Feng and Boyd-Graber, 2019). In our work, participants need to collaborate with explanations to choose the most accurate neural networks among others. Instead of exploring the role of explanations in collaboration, we mainly examine the potential of conversations in aiding users to effectively collaborate with explanations.
(a) Input to the classification model (Swin Transformer)
(b) Explanation generated by Grad-CAM
(c) Explanation generated by LIME
Figure 1. Example explanations generated by Grad-CAM and LIME. (a) is the input to the classification model (Swin Transformer), (b) is the explanation generated by Grad-CAM, and (c) is the explanation generated by LIME. The predicted class of the model is ”Siamese cat”.
3. Methodology
Our study aims to investigate the impact of conversations on the explainability of AI models by observing participants’ comprehension, acceptance, trust of the static explanations, and ability to use the explanations to select the most accurate neural networks before and after the conversation. We devised both objective and subjective evaluations. For the objective evaluation, participants are presented with three neural networks and need to identify the most accurate one based on the provided explanations. All three networks make the same decisions on several images; the only differences between the candidates are the explanations of their predictions. Consequently, to make the correct choice, participants must understand the explanations. Additionally, we gauge participants’ perceptions of static explanations, including understanding, acceptance, and trust. This section delves into a comprehensive overview of our research methodology. Our study has received approval from the Institutional Review Board at the University (#IRB-2023-254).
3.1. Task, AI models and Explanation Method Selection
In our study, we focus on the image classification task — a cornerstone in the field of computer vision (CV) that has been the subject of various human-AI collaborative studies (Taesiri et al., 2022; Jeyakumar et al., 2020). Image classification, by virtue of its minimal domain-specific expertise requirement, lends itself well to crowdsourcing where participants are from diverse domains. We specifically choose the ImageNet dataset (Deng et al., 2009) for this purpose, which boasts 14,197,122 annotations aligned with the WordNet hierarchy and stands as a significant benchmark in image classification and object detection. In our experiments, we select and train three classification models, each exhibiting varied top-1 accuracies: the Swin Transformer (84.1%) (Liu et al., 2021b), VGG-16 (71.6%) (Simonyan and Zisserman, 2015), and AlexNet (56.5%) (Krizhevsky et al., 2012).  In the objective evaluation, participants need to choose, from these three classification models, the best one based on provided explanations (detailed in Section 3.2.1). Alongside the conversation panel in the chat interface (Fig. 3), we display the predictions and the explanations of only the Swin Transformer as a reminder of the task. But participants can still ask questions about any other classification models.
We focus our research on feature attribution explanation methods, which generate explanations by investigating the importance of different input features to final predictions. The reason is that feature attribution explanations are more intuitive to understand for laypeople as they directly show the relationship between input features and model predictions (Kim et al., 2023). Specifically, we select two explanation techniques from two main categories of feature attribution explanation methods: LIME (Ribeiro et al., 2016) (surrogate methods) and Grad-CAM (Selvaraju et al., 2017) (gradient-based methods). LIME works by training an inherently interpretable surrogate model, e.g., a linear model, around a specific prediction. It uses the simple surrogate model to explain the predictions from complex classification models. It spotlights the most crucial features of a model input. Grad-CAM manifests its explanations via a heatmap, color-coding regions of an input image based on their significance to the model’s prediction. It works by calculating the gradients of the output class with respect to the feature maps in the final convolutional layer of a deep neural network and using these gradients to weight the feature maps. The resulting weighted feature maps are then averaged to generate a heatmap that highlights the most activated regions of the image. Figure 2 displays example explanations generated by these two explanation methods. Participants are randomly assigned with an explanation method to do the study once they log into our system.
3.2. Study Design
We devise both subjective and objective evaluations to assess the effect of conversations on participants’ understanding, acceptance, and trust of static explanations. We conduct these evaluations both before and after conversations. By comparing the results, we estimate the impact on conversations for static explanations.
Figure 2. An example of the objective evaluation. The objective evaluation aims to objectively measure participants’ comprehension of static explanations. Each choice contains a prediction from a different classification model, paired with its respective static explanation. Participants need to choose the best model based on the explanations.
Aspect Question
Comprehension
How much do you think you understand the explanations provided for predictions of deep learning models?
Perceived Usefulness
Using explanations would improve my understanding of deep learning models’ predictions.
Using explanations would enhance my effectiveness in understanding predictions of deep learning models.
I would find explanations useful in understanding predictions of deep learning models.
Perceived Ease-of-Use
I become confused when I use the explanation information.
It is easy to use explanation information to understand predictions of deep learning models.
Overall, I would find explanation information easy to use.
Behavioral Intention
I would prefer getting explanation information as long as it is available when getting predictions from deep learning models.
I would recommend others use explanation information to understand predictions of deep learning models.
Trust
How would you rate the competence of the explanation method? - i.e. to what extent does the explanation method perform its function properly?
How would you rate the dependability of the explanation method? - i.e. to what extent can you count on the explanation method to explain predictions of deep learning models?
How would you rate your degree of faith that the explanation method will be able to explain predictions of deep learning models in the future?
How would you rate your overall trust in the explanation method?
Table 1. Detailed questions in the subjective evaluation, which measures participants’ feelings about static explanation from five aspects, i.e., Comprehension, Perceived Usefulness, Perceived Ease-of-Use, Behavioral Intention, and Trust. The user will respond to each question using a 7-point Likert scale.
3.2.1. Objective Evaluation – Decision-Making of Classification Models
The evaluation aims to objectively evaluate participants’ understanding of static explanations. We ask participants to choose, from three classification models, the most accurate on unobserved test data. The three classification models yield identical decisions on 5 images and come with static explanations from the same method. The only differences between the three networks lie in their explanations. Hence, to make the correct selection, the participants must understand the explanations.
We choose three classification models with different performances for this evaluation: Swin Transformer with 84.1% accuracy, VGG-16 with 71.6% accuracy, and AlexNet with 56.5% accuracy. Figure 2 presents a sample image for this evaluation. The full set of images can be accessed in Appendix A. For evaluation integrity, the order of different classification models is random for each participant and the examples used in the evaluation differ from those displayed during conversations.
To ensure that users can make the correct choice based on the explanations, it is crucial that the explanations for model predictions provide clear indicators of each model’s performance. While we recognize that high-performing classification models generally produce more reasonable explanations than low-performing ones, this is not always true for two main reasons. Firstly, the accuracy of a model does not guarantee the plausibility of its explanations. A lower-performing model might align more with human intuition and thus generate seemingly reasonable explanations. Secondly, as we choose examples where all three models make the same predictions, their explanations might be all the same and reasonable. Therefore, we have selected explanation examples where better classification models indeed provide more reasonable explanations. This approach allows users to evaluate and select the best classification models based solely on the explanations, thus making their choice a true reflection of their objective understanding.
An explanation is deemed more reasonable when it focuses on more features that are unique to the predicted class. For example, an explanation that highlights both the head and tail of a fish is considered better than an explanation focusing on the fish head only. This is because a classification model that relies on multiple features for predictions is likely to still make the correct decision if a subset of features are missing or obscured in the image. In the example presented in Figure 2, Model is better than Model A or Model C as Model B utilizes both the head and the body for classification, and does not focus on the background like Model A.
3.2.2. Subjective Evaluation
We also measure participants’ subjective perceptions of static explanations, including their comprehension, acceptance, and trust. Each question is chosen based on an in-depth review of existing literature and has been validated in prior research. Subjective Evaluation contains a total of 13 questions, all of which utilize a 7-point Likert scale for responses. Table 1 lists all questions used in the subjective evaluation.
  • Comprehension (Hoffman et al., 2018; Cheng et al., 2019): Participants’ subjective perceptions of their understanding of explanations. It complements the objective evaluation, providing a holistic perspective on participants’ understanding of static explanations.
  • Perceived Usefulness (Davis, 1989; Davis et al., 1989; Diop et al., 2019): The degree to which participants feel that the explanations enhance their experience with deep learning models. Along with perceived ease of use and behavioral intention, these three aspects measure participants’ acceptance of static explanations. They are derived from the Technology Acceptance Model (TAM) (Davis, 1989; Davis et al., 1989; Diop et al., 2019), a widely applied theory for understanding individual acceptance and usage of information systems. As the explanations are used by end-users, investigating their acceptance of the explanations is very important.
  • Perceived Ease of Use (Davis, 1989; Davis et al., 1989; Diop et al., 2019): Participants’ assessment of the simplicity and clarity of the explanations.
  • Behavioral Intention (Davis, 1989; Davis et al., 1989; Diop et al., 2019): The tendency of participants to utilize the explanation information in the future.
  • Trust (Muir and Moray, 1996): Participants’ confidence in the reliability of the explanation methods to function as intended. Trust has been recognized as an important factor in human-AI collaboration as it mediates the human’s reliance on AI models, thus directly affecting the effectiveness of the human-AI team (Sebo et al., 2020; Seaborn et al., 2021; Doshi-Velez and Kim, 2017). Recent studies show that increased transparency in ML systems doesn’t significantly impact people’s trust (Cheng et al., 2019; Schaffer et al., 2019; Kunkel et al., 2019; Poursabzi-Sangdeh et al., 2021), possibly due to complex trust mechanisms (McKnight et al., 1998; Zhang et al., 2020) and potentially overloaded information (Schaffer et al., 2019; Poursabzi-Sangdeh et al., 2021; Feldhus et al., 2022). In this study, our intention is to discern if clarifying participants’ confusion via conversations about explanations can influence their trust in explanations.
Discipline Number of Participants
Business
8
Psychology
7
Accountancy
6
Biology
6
Economics
5
Chemistry, Chemical Engineering and Biotechnology
4
Sociology
4
Communication Studies
3
Linguistics and Multilingual Studies
3
Civil and Environmental Engineering
2
Computer Science
2
Electrical and Electronics Engineering
2
Sport Science & Management
2
Design (Architecture)
1
Foods science
1
Mathematics
1
Medicine
1
Maritime Studies
1
Physics
1
Table 2. Disciplines of our participants and the number of participants in each discipline. There are 60 participants from 19 different disciplines in total.
3.3. Study Procedure
Before participation, individuals are required to sign an informed consent form that outlines the study’s objectives and procedures. The form also clarifies compensation details and assures both the anonymity and confidentiality of data collected during the study. Upon signing the consent, participants receive an email providing a website link and login credentials (username and password). They are instructed to use these credentials to access the study platform. Details about the study platform are in section 3.6.
After logging in, a pop-up prompt provides an overview of the tasks ahead. Participants then complete pre-conversation objective and subjective evaluations of static explanations. The objective evaluation measures participants’ understanding of static explanations by letting them choose, from three classification models, the most accurate on unobserved test data. The subjective evaluation, with 13 self-reporting questions, probes their perceived comprehension, acceptance, and trust towards the static explanations. Following these evaluations, participants engage in a discussion about static explanations with an XAI expert to deepen their understanding. During the conversation, the example images displayed are different from those used in the evaluations; however, the explanation methods remain the same. Participants are motivated to understand the explanations as they need to select the best-performing classification model using explanations only when doing objective evaluation. Our XAI experts faithfully answer the user’s questions based on their knowledge, trying to help the user gradually understand the explanation. After the conversation, participants complete the same set of evaluations as before. All evaluation outcomes and conversation records are documented. Upon study completion, each participant receives a \$10 reward.
3.4. Wizard-of-Oz Experimental Setting
We conduct conversational explanations with participants in the Wizard-of-Oz (WoZ) setting (Kelley, 1984). Participants interact with a dialogue system that they believe to be autonomous but is actually operated by a human being, the Wizard. The reason for using WoZ is that we can use the collected conversations for future dialogue system development. Several works have shown that the WoZ approach can be applied to collect high-quality typed conversations where a machine learning-based system can learn from(Wen et al., 2017; Budzianowski et al., 2018; El Asri et al., 2017).
During conversations, participants are encouraged to pose questions or offer comments on the provided static explanations. To maintain consistency in the wizard responses, we develop a question-answer repository. We populate an initial repository based on the understanding of the explanation methods (Selvaraju et al., 2017; Ribeiro et al., 2016) and a pilot study with 6 participants. During the actual experiment, when a participant presents a new question, the corresponding answer is added to the question-answer repository. If the same question occurs again, we refer to the previously stored answer. Section 4.5 presents further discussion of the questions from participants during our study.
Figure 3. The Conversation page where users can discuss static explanations with an expert.
3.5. Participant Recruitment and Selection
We recruit our participants in two ways: by posting advertisements on an online forum and by emailing students across various departments and schools. To join our study, participants have to be 21 years old or older, fluent in English, and not involved in research about XAI. To avoid bias in the study and mirror real-world utilization of XAI, we selectively recruit participants to promote the diversity in themajors / discipline of specialization of the participants. Our study is conducted remotely. To ensure data quality, participants are excluded from data analysis if they are away from the keyboard for more than 10 minutes during the study. This results in a total of 60 participants from 19 majors. Table 2 shows statistics of the majors that the participants are enrolled in. Previous research has shown that repeated interactions with AI-based decision-making systems can affect decision subjects’ fairness perceptions (Gemalmaz and Yin, 2022). To mitigate potential bias, each participant can partake in our study only once.
3.6. User Interface design
To conduct the study, we design and build a web-based platform where participants can remotely finish the whole procedure of the experiment. The web-based platform will ensure that all communications between users and XAI experts are text-based and recorded. Our platform will contain three different web pages, i.e., the Login page, the Conversation page, and the Evaluations page. The Login page allows users to log into our platform by using the credentials (username and password) provided to them. After users log onto the platform, they are directed to the Evaluations page. The Evaluations page helps participants complete objective and subjective evaluations (more details in section  3.2.1 and 3.2.2): choosing the most accurate model based on explanations and reporting their perceptions of static explanations. Pre-conversation evaluations are conducted before starting conversations and post-conversation evaluations are conducted after they discuss static explanations with XAI experts. The Conversation page is where participants discuss static explanations with XAI experts. There are two sections on the page. On the left, the user sees a task description, a description of the prediction model, a description of an explanation technique, a model input, a model output, an explanation from the explanation model, and a brief description of the explanation. Within the chatbox on the right, the user can converse with an XAI expert to clarify the explanation. Through a conversation, a user can ask any questions or provide any comments related to the explanation on the left side and our XAI experts will faithfully answer his/her questions based on their knowledge, trying to help the user gradually understand the explanation. Figure 3 shows an example screenshot of the Conversation page.
Explanation Methods Evaluation Timing Objective Understanding (Decision-Making Accuracy) Subjective Understanding Perceived Usefulness Perceived Ease of Use Behavioral Intention Trust
LIME before 0.38 $\pm$ 0.20 4.03 $\pm$ 1.35 5.09 $\pm$ 1.07 4.48 $\pm$ 0.94 5.25 $\pm$ 0.95 4.15 $\pm$ 0.88
after $*{}^{*}$$\pm$ 0.16 $*{}^{*}$$\pm$ 0.88 $*{}^{*}$$\pm$ 0.66 $*{}^{*}$$\pm$ 0.84 $*{}^{*}$$\pm$ 0.81 $*{}^{*}$$\pm$ 0.73
Grad-CAM before 0.82 $\pm$ 0.21 4.17 $\pm$ 0.91 5.49 $\pm$ 0.97 4.71 $\pm$ 0.95 5.52 $\pm$ 0.65 4.40 $\pm$ 1.00
after $*{}^{*}$$\pm$ 0.11 $*{}^{*}$$\pm$ 0.97 $*{}^{*}$$\pm$ 0.60 $*{}^{*}$$\pm$ 0.82 $*{}^{*}$$\pm$ 0.79 $*{}^{*}$$\pm$ 0.80
Table 3. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, perceived usefulness, perceived ease of use, behavioral intention, and trust of different static explanations before and after conversations. Each score is presented as mean $\pm$ standard deviation. $*{}^{*}$ indicates that post-conversation scores are statistically higher than pre-conversation scores with $p<0.001$ using the Student’s t-test.
(a) Results for LIME
(b) Results for Grad-CAM
Figure 4. Results of objective evaluation (decision-making accuracy), participants’ subjective understanding, perceived usefulness, ease of use, behavioral intention, and trust in static explanations, before and after conversational explanations. Decision-making accuracy is ranged from 0 to 1 and the rest scores are from 1 to 7. Figure (a) displays the scores for LIME and Figure (b) displays the average scores for Grad-CAM.
4. Results & Discussion
This section presents the experimental results of the objective and subjective evaluations before and after conversational explanations with XAI experts. We also analyze collected conversations and discuss the implications of our study for building dialogue systems to explain static explanations in the future. We report the statistical significance of the results using Student’s T-tests.
4.1. Conversational explanations improve comprehension of static explanations
[…]
4.2. Objective performance varies between LIME and Grad-CAM
Grad-CAM has a higher accuracy of objective evaluation compared to LIME. A potential reason might be the inherently intuitive nature of the explanations produced by Grad-CAM compared to LIME. This aligns with prior research on different explanations (Kim et al., 2023), where heatmaps were highlighted as more intuitive. Even though Grad-CAM is more intuitive to understand and participants can intuitively choose the best classification model based on the heatmap in the objective evaluation, participants’ self-reporting comprehension score is just slightly larger than 4 (average understanding). This shows that participants still feel confused about how Grad-CAM works and how it explains models’ predictions, even though they can perform well in the objective evaluation.
This also demonstrates that subjective and objective evaluations measure participants’ understanding of static explanations from complementary aspects. Self-reporting scores can be influenced by personal biases, while the objective evaluation might not capture users’ feelings about understanding. Combining both methods can provide a comprehensive view of participants’ understanding of static explanations.
Discipline Evaluation Timing Objective Understanding (Decision-Making Accuracy) Subjective Understanding Perceived Usefulness Perceived Ease of Use Behavioral Intention Trust
Business before 0.63 $\pm$ 0.36 4.00 $\pm$ 0.88 5.26 $\pm$ 1.02 4.43 $\pm$ 1.07 5.29 $\pm$ 0.75 4.32 $\pm$ 0.96
after 0.77 $\pm$ 0.22 5.00 $\pm$ 0.96 5.91 $\pm$ 0.72 5.29 $\pm$ 0.74 5.75 $\pm$ 0.75 5.00 $\pm$ 0.93
Engineering before 0.53 $\pm$ 0.34 4.50 $\pm$ 1.60 5.21 $\pm$ 1.49 4.67 $\pm$ 0.67 5.19 $\pm$ 1.07 4.16 $\pm$ 1.15
after 0.65 $\pm$ 0.26 5.38 $\pm$ 1.19 6.04 $\pm$ 0.57 5.50 $\pm$ 0.76 5.94 $\pm$ 1.15 5.09 $\pm$ 0.79
Humanity pre-conv 0.67 $\pm$ 0.26 4.13 $\pm$ 1.14 5.46 $\pm$ 1.05 4.55 $\pm$ 1.01 5.57 $\pm$ 0.79 4.35 $\pm$ 0.97
post-conv 0.80 $\pm$ 0.22 5.52 $\pm$ 0.85 6.16 $\pm$ 0.63 5.46 $\pm$ 0.98 6.11 $\pm$ 0.74 5.11 $\pm$ 0.78
Science pre-conv 0.51 $\pm$ 0.28 3.93 $\pm$ 1.16 5.09 $\pm$ 0.80 4.78 $\pm$ 0.9 5.30 $\pm$ 0.82 4.18 $\pm$ 0.84
post-conv 0.60 $\pm$ 0.25 5.47 $\pm$ 0.83 5.91 $\pm$ 0.62 5.47 $\pm$ 0.77 5.93 $\pm$ 0.78 5.02 $\pm$ 0.66
Table 4. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, acceptance, and trust of explanation for participants from different disciplines groups before and after conversations. Scores are presented as mean $\pm$ standard deviation. We do not observe statistically significant differences among participants from different disciplines.
4.3. Conversational explanations improve acceptance and trust of static explanations
[…]
(a) Results before conversational explanations
(b) Results after conversational explanations.
Figure 5. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, acceptance, and trust of explanation for participants from different disciplines groups before and after conversational explanations. Decision-making accuracy is ranged from 0 to 1 and the rest self-reported scores are from 1 to 7. Figure (a) shows the results before conversational explanations and (b) shows the results after conversational explanations.
4.4. No statistically significant differences observed among participants from different disciplines
To investigate whether participants from different disciplines perform differently, we categorize their disciplines into the following four groups:
  • Business: Business and Accountancy.
  • Engineering: Civil and Environmental engineering, Computer Science, Electrical and Electronics Engineering, Maritime Studies, and Food Science.
  • Humanities: Psychology, Economics, Communication Studies, Linguistics and Multilingual Studies, and Sociology.
  • Science: Biology, Chemistry, Chemical Engineering and Biotechnology, Sport Science & Management, Mathematics, and Medicine.
Table 4 and Figure 5 display the results of participants from these discipline categories. We do not observe statistically significant differences among different discipline groups. We employ the one-way ANOVA to validate our observations. For pre-conversation results, the significance values are as follows: decision-making accuracy ($p=0.36$), self-reporting comprehension score ($p=0.71$), perceived usefulness score ($p=0.75$), perceived ease of use ($p=0.79$), behavioral intention ($p=0.60$), and trust ($p=0.94$). The post-conversation evaluations show significance values of 0.06, 0.39, 0.58, 0.91, 0.63, and 0.97, respectively.
Participants from different disciplines do not seem to benefit from discipline-specific training when trying to understand the explanations. Although one may be tempted to conjecture that people from different disciplines would possess different skills, which may lead to easier or more difficult experiences working with AI, our results do not reveal any such differences. This may suggest that understanding AI and XAI requires highly specialized background knowledge that are not shared by the four broad categories we used in the study. Hence, the duty falls upon the XAI designers to effectively communicate with the users wihout assuming any knowledge background. Conversational explanations may function as one such form of communication.
Question Category
Question Examples
Num
Basic concepts in machine learning
  • $\bullet$
    What is a deep learning model?
  • $\bullet$
    What is the image classification task?
  • $\bullet$
    How does the model know what features to extract?
85
Application, performance, and limitations of machine learning models
  • $\bullet$
    How about the precision of the classification model?
  • $\bullet$
    Where has this swin Transformer classification method been used in practical applications?
  • $\bullet$
    Will the different species of an animal affect the classification model categorizing the animal?
68
Diagram reading
  • $\bullet$
    Are regions colored in red areas that have been identified as containing key features for the animal?
  • $\bullet$
    What are the yellow line spots for (in LIME explanations)?
  • $\bullet$
    What do the red color and blue color mean (in Grad-CAM explanations)?
16
Basic concepts of explanation methods
  • $\bullet$
    What is the explanation model used for?
  • $\bullet$
    Can LIME be used without the internet?
  • $\bullet$
    What are some limitations of the Grad-CAM (LIME) method?
95
Mechanism of explanation methods
  • $\bullet$
    Why does the (LIME) explanation not highlight all the parts of the leopard?
  • $\bullet$
    How LIME model recognize the most important parts for the model prediction?
  • $\bullet$
    Seems like the Classification Model and the Explanation Model are trained separately - how can we be sure that the underlying logic of making a prediction is the same for both models?
91
New explanations
  • $\bullet$
    Can you list other visualization methods?
  • $\bullet$
    Is there anything special about the Grad-CAM (or LIME) method that is different from others?
  • $\bullet$
    What if there are both fishes and humans in an image? How should this image be classified, and can you provide such explanations?
42
Table 5. Overview of Collected Questions. Including categories of questions, examples, and the count of questions in each category. Questions are collected based on conversations between participants and XAI experts about an explanation generated by Grad-CAM or LIME.
4.5. Analysis of Collected Conversations
We collect 60 free-form conversations between XAI experts and participants from 19 different disciplines. On average, each conversation had 27.4 turns, with each turn comprising approximately 14.4 tokens. By analyzing the users’ questions, we divide them into four categories:
  • Basic concepts in machine learning: Questions about basic terms and concepts in machine learning that lay people may not know, e.g., what is a deep learning model, what is accuracy, the model structure, and the training data, etc.
  • Application and performance of machine learning models: Questions about the ability, accuracy, and limitations of machine learning models.
  • Diagram reading: Questions about the explanation diagram generated by Grad-CAM or LIME, e.g., what different color represents in the heatmap.
  • Basic concepts in explainable AI: Questions about basic concepts of explanation methods, e.g., what are explanation methods?
  • Mechanism of explanation methods: Questions about how explanation methods work and how the provided explanation is generated.
  • New explanations: Questions that require the generation of new explanations, including from other explanation methods, comparison between the provided explanation and other, currently unavailable explanation methods, or explanations for different predictions.
Based on this categorization, we build a repository for questions that could occur in the conversations. In total, we collect 397 questions from the four different categories. Table 5 contains examples and the number of questions in each category. As observed in Table  5, the questions of participants mainly revolve around basic concepts in machine learning, the fundamentals of explanation methods, and their underlying mechanisms. This trend might be attributed to the multi-disciplinary of the participants. It suggests that many participants may not be familiar with machine learning models and explanation methods, which is aligned with the real application of explanation methods. Therefore, it’s crucial to tailor responses to these questions to help users better understand explanations. Furthermore, we note a marked interest in new explanations. This could indicate that as users become more familiar with provided explanation examples, they exhibit curiosity about alternative explanation methods and how models might behave under specific scenarios. Concurrently, the diagram reading category contains only 16 questions, implying that explanations generated by Grad-CAM and LIME were relatively straightforward and easy to understand. The diverse range of questions sourced from our conversations underscores that static, one-off explanations are often insufficient for users to understand them. Engaging in dialogue can provide more dynamic and tailored explanations to users, hence deepening their understanding of static explanations.
4.6. User questions are unanticipated and diverse
Having well internalized their knowledge, experts are often unable to estimate what laypeople know (Wittwer et al., 2008). This phenomenon is also referred to as the “curse of knowledge”(Camerer et al., 1989). As a result, experts tend to overlook potential areas of confusion or make unwarranted assumptions about what is “common knowledge”. While analyzing the collected conversations, we often find ourselves unable to anticipate the user questions, which corroborates the literature. We describe a few examples below.
Several participants misunderstand the idea of the heatmap produced by Grad-CAM as depicting literal heat dissipating from objects. They infer that the model uses the temperature of objects to perform classification. In reality, a heatmap is just a metaphor that visualizes numerical values distributed spatially, which refer to the feature importance in our case. This misconception leads to questions about how the heat of objects is measured and why non-living objects are warmer than their environment. Some example utterances from participants include: “So the Grad-cam method basically just refers to the usage of generating a heatmap to capture living matters correct? … based on the parts of the image that generate more heat?” –P36, “basically using heat to predict what is the input right?…how will we know what is the animal or input simply based on heat?” – P47, “if these are pictures, how do they figure out the heat since the animal isn’t generating heat” – P49, “So a heat sensor is not required? A heatmap is automatically generated from each photo and analyzed using the model.” – P52.
A second common misconception is the conflation between the explanation technique and the classification models. Some example user questions include: “is the explanation method what the model uses to classify & predict what the image is supposed to be?” – P6, “swin transformer uses LIME model? … what’s the differences between lime model and swin transformer?” – P8. Furthermore, participants face challenges in understanding certain terms commonly used in AI and XAI, even though these terms are frequently used and understood within academic communities. Many participants asked questions about basic concepts in machine learning, such as: “what is explanation method?” – P7, “how do you classify the image?” – P17, “what is the algorithm? does it means lime? what are deep neural networks?” – P32, “How would you explain the term ”perturbations of images” to a five-year-old?” – P46.
The observations from the interactions between XAI experts and layperson users demonstrate the importance of conversations for users to understand static explanations as they bridge the knowledge gap between the two groups. Conversations can reveal the specific areas of misunderstanding, assumptions users might make, and knowledge they might lack. Hence, conversational explanations may help the AI system communicate with and bring genuine understanding to the users.
4.7. Implications for building dialogue systems to explain static explanations
Our study indicates the impact of conversational explanations on user comprehension, acceptance, and trust of static explanations. Static explanations, while informative, may not cater to users with varied backgrounds and expertise. Engaging in conversational explanations provides a dynamic and interactive medium for users to seek clarifications, ask questions, and thereby facilitate a deeper and more personalized understanding.
The emergence of advanced conversational agents (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question-answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023) paves the way to build dialogue systems that can explain model decisions and explanations in conversations. Below, we explore the implications of our study for designing such dialogue systems.
  • Extensive Coverage of AI and XAI Concepts. As observed in our study, there are a large portion of questions are the core concepts of machine learning models and explanation methods. To answer those questions, dialogue systems need to be trained on a comprehensive corpus encompassing AI and XAI concepts. Besides, in our study, participants also are curious about the applications, performances, and limitations of machine learning models and explanation methods. Therefore, besides answering theoretical questions, dialogue systems also should be able to link them to real-world applications and limitations.
  • Capability to Generate New Explanations. As an improved understanding of the provided explanations, participants in our study exhibit curiosity about alternative explanation methods and different explanation examples. Dialogue systems should be designed to provide new explanations to users. They should not only retrieve stored explanations but also construct new ones based on the evolving understanding of the user’s question. For instance, if a user is curious about how changing a feature would affect a model’s output, the system should dynamically generate a new explanation, elucidating the potential impact.
  • Ability to Interpret Diagrams. A significant portion of AI and XAI explanations often comes in the form of diagrams (Selvaraju et al., 2017; Ribeiro et al., 2016), such as heatmaps from Grad-CAM or feature importance via LIME. While our study observed fewer questions about diagrams, those questions are essential for users to understand explanations. Therefore, future dialogue systems should have visual processing capabilities, understanding and interpreting diagrams contextually. For instance, they should be able to recognize colors, patterns, and other graphical elements in heatmaps or charts and relate them to users’ questions. The recent development in multimodal large language models (Zhu et al., 2023; Driess et al., 2023; Gong et al., 2023) is a promising direction to achieve this goal.
5. Limitation and Future Work
Our study exploits the effects of free-form conversations that accompany and clarify static explanations. Despite the insights gained, there are several limitations that should be acknowledged. First, the static explanations used in our study are limited. Our experiments focused on feature attribution explanation methods; we selected LIME and Grad-CAM, which are representative of gradient-based methods and surrogate methods respectively. The applicability of our findings to other explanation methods, such as example-based explanation methods, remains an open question. Second, as our main objective was to discern the effects of free-form conversational explanations, we did not delve into the comparative performance of different explanation methods. In our experiments, we intentionally selected explanation examples where the best classification model yielded more reasonable explanations. The explanation examples discussed by participants and XAI experts were chosen such that they reasonably explain the predictions of the classification model. Future work would be to extend these conversations to include explanations that might be less reliable. Third, we investigate how conversations can affect user trust in explanations of AI systems. It’s noteworthy that prior studies have highlighted that humans may place undue trust in AI models even when system decisions are flawed (Wang and Yin, 2021; Zhang et al., 2020). Within the scope of the current study, we do not specifically evaluate if this trust exhibited by users is misplaced or justified. Fourth, we explore how conversations foster user trust in explanations in our study. Nevertheless, previous studies (Wang and Yin, 2021; Zhang et al., 2020) have shown that humans may trust AI models even if they make wrong decisions. We do not explore whether users’ trust in our study is misplaced trust or not, which can be extended in the future. Finally, our research is confined to one (anonymized) geographical region. Cultural nuances may influence the behaviors of users of XAI and how they seek to clarify confusion. Future studies could involve recruiting participants from diverse countries or regions to draw a more comprehensive conclusion.
6. Conclusion
In our work, we conduct Wizard-of-Oz experiments to investigate how free-form conversations assist users in understanding static explanations, promoting trust, and making informed decisions about AI models. Participants engage in conversational explanations with XAI experts with the aim of understanding how the provided static explanation explains the model decision. To evaluate the effects of conversations, we design objective and subjective measurements. We observe a notable improvement in users’ comprehension, acceptance, trust, and collaboration after conversations. From collected conversations, we find that participants’ questions and confusions are diverse and unanticipated. Our findings advocate for the integration of dialogue systems in future XAI designs to ensure more personalized explanations. We hope that our study design and insights will aid the development of conversational XAI research.
Appendix A Objective Evaluation – Decision-Making of classification models.
The evaluation aims to objectively evaluate participants’ understanding of static explanations. We ask participants to choose, from three classification models, the most accurate on unobserved test data. All three classification models make the same decisions on 5 images, accompanied by static explanations from the same explanation method. The only differences between the three networks lie in their explanations. Hence, to make the correct selection, the participants must understand the explanations.
Figure 8 presents the full set of images listed in the objective evaluation for Grad-CAM, while Figure 11 showcases the same for LIME.
Figure 6. Objective evaluation questions used for Grad-CAM.
Figure 7. Objective evaluation questions used for Grad-CAM.
Figure 8. Objective evaluation questions used for Grad-CAM.
Figure 9. Objective evaluation questions used for LIME.
Figure 10. Objective evaluation questions used for LIME.
Figure 11. Objective evaluation questions used for LIME.